The cost of the downloads in$ is c=2v,where c is the total cost of the downloads and v is the number of videos. The constant of proportionality in the equation in terms of cost per video is? In this example,the constant of proportionality is also the ? Of the videos
Found 2 solutions by addingup, macston:
Answer by addingup(3677)

The cost C is variable, it changes with the change in V, the number of videos you rented. The constant is 2, this is the cost per video regardless of number of videos you rent. So you'd have
C= 2V
2= 2(1)
4= 2(2)
6= 2(3)
8= 2(4)
and so on, where your cost goes up by 2 for every additional video you rent.
J

Answer by macston(5194)

.
The constant of proportionality in the equation in terms of cost per video is?
The only constant in the equation is 2.
.
In this example, this constant is also the cost per video ( in dollars) to download.